### v1

# DBellQuant: Breaking the Bell with Double-Bell Transformation for LLMs Post Training Binarization

Abstract
Large language models (LLMs) demonstrate remarkable performance but face substantial computational and memory challenges that limit their practical deployment. Quantization has emerged as a promising solution; however, its effectiveness is often limited by quantization errors arising from weight distributions that are not quantization-friendly and the presence of activation outliers. To address these challenges, we introduce DBellQuant, an innovative post-training quantization (PTQ) framework that achieves nearly 1-bit weight compression and 6-bit activation quantization with minimal performance degradation. DBellQuant uses Learnable Transformation for Dual-Bell (LTDB) algorithm, which transforms single-bell weight distributions into dual-bell forms to reduce binarization errors and applies inverse transformations to smooth activations. DBellQuant sets a new state-of-the-art by preserving superior model performance under aggressive weight and activation quantization. For example, on the Wikitext2 dataset, DBellQuant achieves a perplexity of 14.39 on LLaMA2-13B with 6-bit activation quantization, significantly outperforming BiLLM’s 21.35 without activation quantization, underscoring its potential in compressing LLMs for real-world applications.
1 Introduction
In recent years, the rapid advancement of large language models (LLMs) demonstrate exceptional performance in a variety of complex tasks that involve natural language understanding and generation achiam2023gpt ; dubey2024llama . However, these models often comprise hundreds of billions of parameters, posing significant challenges for their deployment in real-world applications because of the substantial computational and memory requirements (e.g. a 70B model requires around 150GB GPU memory), resulting in huge operational costs and unacceptable inference latency.
In this context, quantization, as an effective model compression technique, has garnered significant attention. Due to the sparsity of information density in the weights of LLMs wang2024bitnet ; yu2024super , weight binarization has emerged as a promising quantization scheme. Methods such as PB-LLM shang2023pb and BiLLM huang2024billm have significantly reduced the errors introduced by binarization by applying finer-grained processing to critical weights. Despite its theoretical advantages, the application of quantization in compressing LLMs continues to face numerous challenges egashira2024exploiting . For instance, systematic outliers in activation values can lead to substantial quantization errors, negatively impacting model accuracy sun2024massive . Although recent studies xiao2023smoothquant ; shao2023omniquant have proposed mitigating such errors by redistributing the scaling factors between weights and activation values. Other research directions ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm focus on leveraging Hadamard transformations to effectively eliminate outlier features. However, no binarization methods have taken care of simultaneous smoothing activation outliers. Consequently, the demand of higher bit-width activations during computation leads to significant computational overhead. BitNet a4.8 wang2024bitnet addresses this issue by employing quantization-aware training (QAT) to achieve a 1-bit LLM with 4-bit activations. Nevertheless, the QAT approach requires extensive computational resources and prolonged training time. Therefore, developing post-training quantization (PTQ) methods for binary compression that minimize performance loss while simultaneously smoothing activation and thus reducing activation bit-width remains a critical and challenging task.
Figure 1: Performance on Wikitext2 dataset. DBellQunat outperforms weight-only quantization method under 8-bit activation setting.
To this end, we begin by revisiting the distribution characteristics of activations and weights in LLMs (Fig. 2(a)). We observe that the unimodal nature of weight distributions leads to substantial quantization errors, particularly in the case of low 1-bit quantization. Ideally, a dual-bell-shaped weight distribution (Fig. 2(b)) can effectively reduce binarization errors. This motivates a key question: Is it possible to transform weights into a dual-bell shape distribution while simultaneously addressing activation outliers to facilitate both activation quantization and weight binarization?
Building on these observations, we propose DBellQuant, a novel weight-activation quantization framework for efficient post-training quantization (PTQ). DBellQuant enables activation quantization while achieving near 1-bit weight compression with minimal accuracy loss. At the core of DBellQuant is a learnable transformation matrix that maps the weight distribution into a dual-bell form, while its inverse is applied to the input activations to smooth them. We begin by analyzing the characteristics of weight distributions suitable to binarization and theoretically derive the feasibility of applying a dual-bell transformation. Based on this, we then develop an efficient and lightweight algorithm, Learnable Transformation for Dual-Bell (LTDB). LTDB initializes the transformation matrix using an activation-aware strategy and optimizes it via a custom objective function that encourages the weights to cluster around two centers. This drives the formation of a symmetric dual-peak distribution, and an early stopping strategy is employed to ensure stable and efficient optimization. Through this equivalence-preserving transformation, weights originally exhibiting unimodal distributions– challenging for binarization– are mapped into near-symmetric dual-bell distributions (Fig. 2(b)), significantly reducing binarization error. Simultaneously, the inverse of the learned transformation is applied to the input activations, effectively scaling down outliers and smoothing the distribution. This makes the activations more amenable to quantization without altering the model’s output (Fig. 2(b)).
Experimental results demonstrate that the equivalent transformation strategy of DBellQuant significantly reduces the loss caused by weight binarization. For the first time under PTQ conditions, it achieves near 1-bit weight compression while simultaneously compressing activations to 6 bits.Across various LLMs and evaluation metrics, DBellQuant consistently achieves state-of-the-art results. For instance, on the Wikitext2 benchmark, we achieves a perplexity of 14.39 on LLaMA2-13B using only 6-bit activations (Fig. 1), significantly surpassing the performance of BiLLM, a method that only quantizes weights, which achieves a perplexity of 21.35.
Figure 2: (a) Before applying DBellQuant, activations exhibit significant outliers, making quantization challenging, while the single-bell-shaped weight distribution hinders binarization. (b) After applying DBellQuant, activations are smoothed with substantially fewer outliers, facilitating easier quantization. Weight distribution is transformed to dual-bell form, which is more conducive to binarization.
2 Related Work
Quantization for Large Language Models
The massive parameter size of LLMs poses significant challenges in terms of memory consumption and computational efficiency. Therefore, quantization is crucial to compress these models, reducing resource requirements while preserving performance for practical deployment. LLMs quantization have introduced a variety of innovative techniques to enhance efficiency while maintaining accuracy. Works like GPTQ frantar2022gptq and OBQ frantar2023optimalbraincompressionframework minimizes reconstruction error by adjusting the remaining unquantized parameters in the block to compensate for the accuracy loss caused by quantization. LLM.int8()dettmers2022gpt3 and ZeroQuantyao2022zeroquant improve quantization accuracy by introducing additional grouping labels for customized quantization blocks. Other works like SmoothQuantxiao2023smoothquant and OmniQuant shao2023omniquant addresses activation outliers by redistributing scaling factors between weights and activations, migrating the quantization difficulty from activation to weights. Additionally, recent approaches leverage Hadamard transformations to suppress activation outliers ashkboos2024quarot ; liu2024spinquant ; sun2025flatquantflatnessmattersllm , while incoherence processing has been proposed for effective low-bit quantization chee2024quip2bitquantizationlarge ; tseng2024quipbetterllmquantization . Collectively, these advancements demonstrate that quantization techniques can be successfully scaled to multi-billion-parameter models, achieving substantial reductions in memory consumption and inference latency without compromising model performance.
Binary Quantization
Binary quantization, an extreme low-bit quantization technique that reduces model weights and activations to binary values (e.g., -1 and +1 or 0 and 1), has gained significant attention for its ability to drastically cut memory usage and computational complexity, making it ideal for resource-constrained devices and efficient deployment of large-scale models. However, applying binary quantization to LLMs presents substantial challenges due to their sensitivity to precision loss, particularly in attention mechanisms and large embedding layers. BinaryBERT bai2020binarybert explored binary quantization for BERT, proposing selective preservation of critical weights in higher precision to mitigate performance degradation. In another direction, PB-LLM shang2023pb introduced a partially-binarized approach for LLMs, retaining a small fraction of salient weights in higher precision while binarizing the rest, enabling extreme low-bit quantization without sacrificing linguistic reasoning capabilities. Recent advancements include structural binarization techniques that leverage novel sparsity forms and standardized importance metrics to selectively binarize and sparsify LLM weights dong2024stbllmbreaking1bitbarrier , as well as strategies like alternating refined binarization and column-group bitmap methods to effectively reduce quantization error and address column deviations li2024arbllmalternatingrefinedbinarizations . These innovations collectively advance the feasibility of binary quantization for LLMs, pushing the boundaries of efficiency without compromising performance.
Figure 3: DBellQuant Framework Overview: (a)First, we can see that the origin weight distribution is single-bell. (b)We utilize Activation-aware initialization to generate origin transformation matrix. (c)We employ the LTDB algorithm for iterative training of the transformation matrix, applying the proposed Dual-Transformation Loss in two ways: for training and as the termination criterion for the training process. (d)The weight distribution after transformation will be double-bell.
3 Method
We begin by exploring the process of binarization, analyzing and theoretically proving the weight distributions suitable for binarization in Sec. 3.1. Based on our analysis, we propose the Learnable Transformation for Dual-Bell (LTDB) algorithm in Sec. 3.2. After investigating the potential of utilizing a learnable transformation matrix to achieve the objective, we redesigned an efficient learnable transformation along with a reasonable activation-aware initialization method, taking into account training difficulty and task complexity. An overview of the algorithm is included in Fig. 3.
3.1 Binarization-Friendly Weight Redistribution
By utilizing the sign function, binarization can convert weights in LLMs into binary values. The per-channel binarization and de-binarization process is as follows:
$\small\beta=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{W_{i,j}},\quad\boldsymbol{\widetilde{W}}=\text{Sign}(\boldsymbol{W}-\beta),\quad\alpha=\frac{1}{n}\sum_{i=1}^{n}|\boldsymbol{W_{i,j}}-\beta_{j}|$ (1)
$\text{Sign}(\boldsymbol{W_{i,j}})=\begin{cases}+1,&\text{if }\boldsymbol{W_{i,j}}>0,\\ -1,&\text{if }\boldsymbol{W_{i,j}}\leq 0,\end{cases},\quad\boldsymbol{W_{deq}}=\boldsymbol{\widetilde{W}}\cdot\alpha+\beta$ (2)
where $\beta$ is the shifting factor and $\alpha$ is the scaling factor for binarization. Previous studies huang2024slim have shown that neural network weights exhibit structured distributions along the channel dimension, with certain channels being more salient. The overall weight distribution typically follows a quasi-Gaussian pattern dettmers2022llmint88bitmatrixmultiplication , as does the channel-wise distribution (Fig. 6). Binarizing such weight matrices introduces significant quantization errors, which can severely degrade model performance.
In LLM binarization, a dual-bell distribution is theoretically more advantageous than a single-bell distribution due to its natural separation into two distinct clusters, which aligns well with binary quantization levels (e.g., -1 and 1), thereby minimizing quantization error. In contrast, single-bell distributions, concentrated around a single peak, often cause significant overlap when mapped to binary values, reducing representation accuracy (see Appendix A.4 for detailed analysis). However, LLM weight distributions typically exhibit single-bell characteristics, and conventional PTQ methods fail to effectively transform them for binarization. While QAT can reshape weight distributions into a dual-bell form through its learning objectives wang2023bitnetscaling1bittransformers , it requires substantial computational resources and prolonged training. To address this, we propose a more efficient PTQ method that rapidly converts single-bell distributions into dual-bell ones, optimizing binary quantization without the need for resource-intensive retraining.
3.2 Learnable Transformation for Dual-Bell Quantization
Learnable Transformation with Auxilary Matrix
As mentioned before, double-bell distributions are advantageous for binarization. However, the key question lies in how to transform a weight matrix that originally follows a single-bell distribution into a double-bell one and ensures the computational results remain unchanged. In this section, we first explore the feasibility of achieving such a transformation through the application of an auxiliary matrix:
Theorem 1.
Let $\boldsymbol{W}\in\mathbb{R}^{n\times m}$ be a weight matrix where each channel $\boldsymbol{w}_{i}$ (for $i\in\{1,2,\dots,n\}$) is sampled from a single-bell Gaussian distribution $\boldsymbol{w}_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. There exists a learnable matrix $T\in\mathbb{R}^{m\times m}$, such that the channels of the transformed matrix $\boldsymbol{W}^{\prime}=\boldsymbol{W}T$ follow a double-bell distribution, specifically a mixture of two Gaussians:
$\boldsymbol{w}_{i}^{\prime}\sim\pi\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(\mu_{2},\sigma_{2}^{2}),$
where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are parameters of the doubel-bell distribution. More detailed proof is shown in Appendix. A.3.
Theorem 1 demonstrates that transforming weight distributions from single-bell to double-bell can be achieved by introducing an auxiliary matrix $T$. However, this approach presents several significant challenges. First, in LLMs, weight matrices typically have extremely high dimensionalities, such as $(4096,4096)$, meaning that the auxiliary matrix $T$ would also be of similarly large dimensions, making it computationally expensive and difficult to learn. Second, to maintain computational consistency, it is necessary to simultaneously apply $T^{-1}$ to the activations, which raises a critical issue regarding the invertibility of $T$. Ensuring strict invertibility introduces additional constraints and complex design steps, further complicating the process. Third, even if $T$ is strictly invertible, it remains uncertain whether this design effectively facilitates activation quantization, as there are no explicit mechanisms in the current approach to optimize activation quantization. These limitations highlight the need for a more efficient and robust design to address the computational and practical challenges associated with auxiliary matrix-based transformations.
Learnable Equivalent Transformation
To address these challenges, we propose a simpler and more efficient method for achieving the transformation. In this approach, the matrix $T\in\mathbb{R}^{1\times C_{\text{in}}}$ to be learned is reduced to a $1\times C_{\text{in}}$ matrix. Compared to the matrix introduced above, this significantly reduces the dimensionality of $T$ to compared to its original size, making it substantially easier to learn.
Furthermore, this approach allows for straightforward transformations to ensure computational consistency without introducing additional complexity as follows:
$\small\textbf{Y}=\textbf{X}*\boldsymbol{W}=\textbf{X}*(T^{-1}*T)*\boldsymbol{W}=\left(\textbf{X}\odot T^{-1}\right)*\left(T\odot\boldsymbol{W}\right)$ (3)
where $X\in\mathbb{R}^{N\times C_{\text{in}}}$ is the input matrix, $N$ is the token length and $C_{\text{in}}$ is the input channel size. $w\in\mathbb{R}^{C_{\text{in}}\times C_{\text{out}}}$ is the weight matrix, where $C_{\text{out}}$ is the output channel size. $\odot$ denotes elementwise multiplication.Moreover, this equivalent transformation matrix $T$ will be directly fused into the LayerNorm weights and the corresponding linear weights, without introducing any additional parameters. However, directly solving this matrix is highly challenging. To address this, we propose a learnable approach to train and derive the matrix effectively.
Here we introduce the way to initialize the learnable transformation matrix $T$ using the following equation:
$\small T_{j}=\frac{\max(|\mathbf{X}_{j}|)^{\epsilon}}{\max(|\boldsymbol{W}_{j}|)^{1-\epsilon}}$ (4)
where $\epsilon$ is a hyperparameter. This initialization strategy provides significant advantages for both weight binarization and activation smoothing. For weight quantization, specifically, when $\max(|\boldsymbol{W}_{j}|)$ is particularly small, it indicates that the absolute value of weights are relatively small, resulting in a large value of $\frac{1}{\max(|\boldsymbol{W}_{j}|)}$ which corresponds to scaling up these smaller weights. Conversely, when $\max(|\boldsymbol{W}_{j}|)$ is particularly large, it reflects larger absolute value weights, which lead to a smaller value of $\frac{1}{\max(|\boldsymbol{W}_{j}|)}$, which will scale down the larger weights. All values can be shifted closer to two central points through these two processes, and it will reduce the quantization error shown in Appendix. A.4. Regarding activation quantization, the initialization explicitly accounts for outliers in the activation matrix, making it inherently activation-aware. This ensures that even without further optimization of activation quantization during subsequent training, it supports near 1-bit weight quantization while effectively reducing activations to a low bit-width.
3.3 Dual-Transformation Optimizing Objectives
Dual-Target Minimum Deviation Loss
Our learning objective is to encourage all weight values to move closer to the two mean centers calculated by Eq. 2, ultimately forming a doubel-bell distribution with these two values as its peaks. During the binarization process, we denote these two points as $m_{1}$ and $m_{2}$ respectively. The simple way to set loss function is as follows:
$\small\mathcal{L}_{\text{DTMD}}=\frac{\lambda_{\text{DTMD}}}{n}\sum_{i=1}^{n}\min\left(\left|\boldsymbol{W}*{T_{i}}-m_{1,i}\right|,\left|\boldsymbol{W}*{T_{i}}-m_{2,i}\right|\right)$ (5)
where $\lambda_{\text{DTMD}}$ represents the coefficient of $\mathcal{L}_{\text{DTMD}}$. However, employing this type of loss function to train the transformation matrix introduces an issue. Specifically, we observed that the transformation matrix tends to shrink progressively during training, contrary to the intended effect of scaling up the originally smaller absolute values. This unintended behavior results in a significant problem: it effectively shifts the quantization challenge from the weights to the activations.
Dual-Target Normalized Proportional Loss
As discussed before, DTMD alone is not enough to train a better weights distribution for binarization. So we introduce a new loss function as follows:
$\small\mathcal{L}_{\text{DTNP}}=\frac{\lambda_{DTNP}}{n}\sum_{i=1}^{n}\begin{cases}\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|}{\left|\text{m}_{1,i}\right|},&\text{if }\left|\boldsymbol{W}*{T_{i}}-\text{m}_{1,i}\right|<\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|\\[10.0pt] \frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{2,i}\right|}{\left|\text{m}_{2,i}\right|},&\text{otherwise.}\end{cases}$ (6)
Figure 4: Dual-Target Minimum Deviation Loss value over iterations across different layers.
where $\lambda_{DTNP}$ represents the coefficient of $\text{Loss}_{\text{rel}}$. By leveraging the dual-target normalized proportional objective, the transformation matrix can be effectively trained to meet the desired behavior, scaling down larger absolute values and scaling up smaller absolute values to approach a double-bell-shaped distribution. Since our target values have been transformed into $\frac{\left|\boldsymbol{W}*{T_{i}}-\text{m}_{i}\right|}{\left|\text{m}_{i}\right|}$ , the final convergence values might not align with our desired results. Furthermore, we propose an early stopping mechanism to prevent the function from converging to an undesired solution that deviates from our intended objective. We observe that by using this loss to train, DTMD drops quickly first and then slowly grow as shown in Fig. 4. Therefore, we introduce an early stop mechanism and DTMD is utilized as a condition for stopping the training.
3.4 Impact of the Inverse of Learnable Transformation Matrix on Activation Smoothing
Figure 5: Visualization of distribution of values in $T$ and $T^{-1}$ of Llama2-7B.
Through the use of DTNP for training, we understand that the transformation matrix $T$ drives all values in the weights closer to the two mean centers calculated by Eq. 2. Prior research has shown that the distribution of weights tends to approximate a quasi-Gaussian distribution, with the majority of values being extremely small and close to zero, while only a very small fraction exhibit relatively large absolute values dettmers2022llmint88bitmatrixmultiplication . Theoretically, this implies that $T$ will contain many values greater than 1 to amplify the numerous near-zero absolute values, bringing them closer to the mean centers. At the same time, very few values of $T$ will be less than 1 to reduce the relatively rare large absolute values, aligning them similarly with the mean centers. In fact, when we visualize $T$ after training as shown in Fig. 5, we observe that less than 5% of its values are below 1. Consequently, the corresponding $T^{-1}$, which is multiplied with activations, has over 95% of its values below 1. This significantly reduces the quantization range of activations and suppresses the magnitude of outlier values within the activations. Visualization results can be seen in Appendix. A.6. As a result, this approach is particularly effective in further facilitating activation quantization.
3.5 Algorithm
The process of Learnable Transformation for Dual-Bell Transformation Algorithm is shown in Algorithm. 1, which adjusts the full-precision weight matrix $W$ using an activation-aware initialization transformation matrix $T$ over $N$ epochs. The algorithm iteratively minimizes dual-target loss functions to guide $W$ toward a dual-bell distribution while employing an early stopping mechanism based on the DTMD.
Algorithm 1 Learnable Transformation for Dual-Bell Transformation (LTDB)
1:function LTDB($\boldsymbol{W},\mathbf{T},N$)
2:     Input: $\boldsymbol{W}\in\mathbb{R}^{n\times m}$ a full-precision weight matrix.
3:     Input: $\mathbf{T}\in\mathbb{R}^{1\times m}$ an activation-aware initialization transformation matrix.
4:     Input: $N$ the total number of epochs.
5:     Output: $\widetilde{\boldsymbol{W}}\in\mathbb{R}^{n\times m}$ the transformed weight matrix.
6:     for $\text{iter}=1,2,\dots,N$ do
7:         $\widetilde{\boldsymbol{W}}\leftarrow\mathbf{T}\odot\boldsymbol{W}$triangleright Perform element-wise multiplication
8:         $\mathcal{L}_{\text{DTMD}},\mathcal{L}_{\text{DTNP}}\leftarrow\text{LossFunc}(\widetilde{\boldsymbol{W}})$triangleright Compute the two dual-target losses for $\widetilde{\boldsymbol{W}}$
9:         $\mathcal{L}_{\text{DTNP}}.backward()$triangleright Use Dual-Target Normalized Proportional Loss for training
10:         if $\mathcal{L}_{\text{DTMD}}>\mathcal{L}_{\text{DTMD-Minimum}}$ then
11:              break $\triangleright$ Stop training based on Dual-Target Minimum Deviation Loss
12:         end if
13:         $\mathcal{L}_{\text{DTMD-Minimum}}\leftarrow\mathcal{L}_{\text{DTMD}}$
14:     end for
15:     return $\widetilde{\boldsymbol{W}}$
16:end function
4 Experiments
4.1 Settings
All experimental procedures were executed utilizing the PyTorch paszke2019huggingface framework in conjunction with the Huggingface library paszke2019huggingface . Models with parameters smaller than 8B are running on a single NVIDIA A30 GPU equipped with 24GB of memory, others are running on a single NVIDIA A100 GPU equipped with 80GB of memory. Consistent with methodologies outlined by Frantar et al. frantar2023gptq and Huang et al. huang2024billm , a calibration dataset comprising 128 samples sourced from the C4 collection raffel2020exploring was employed.
Models and Datasets
Comprehensive evaluations were carried out across several large language model families, including LLaMA, LLaMA-2, and LLaMA-3 touvron2023llama and the OPT series zhang2022opt . The efficacy of the developed DBellQuant was assessed by calculating the perplexity of the models’ generated text on standard benchmarks: WikiText2 merity2017pointer , and a subset of the C4 data raffel2020exploring . Furthermore, the models’ performance was evaluated based on accuracy across seven zero-shot question-answering tasks: ARC-c clark2018think , ARC-e clark2018think , Hellaswag zellers2019hellaswag , PIQA bisk2020piqa , and Winogrande sakaguchi2020winogrande .
Comparison Methods
The primary benchmark for comparison for our DBellQuant approach is BiLLM huang2024billm , which represents the current baseline PTQ technique for binary large language models. Additionally, we include other contemporary PTQ algorithms in our comparison, namely Round-to-Nearest (RTN), GPTQ frantar2023gptq , PB-LLM shang2023pb and the current state-of-the-art PTQ technique ARB-LLM li2024arbllmalternatingrefinedbinarizations .
[…]
Method Activation Bits OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-1-7B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
Full Precision 16 14.62 12.47 10.86 5.68 05.47 04.88 3.32
RTN 16 17165.72 36516.69 11550.91 168388.00 157058.34 47902.32 160389.91
GPTQ 16 14844.73 14114.58 10622.81 267001.72 115905.67 9387.80 14219.35
PB-LLM 16 265.52 124.35 105.16 102.36 69.20 151.09 28.37
BiLLM 16 69.97 49.55 35.36 35.04 32.48 21.35 13.32
[…]
BiLLM 8 88.95 68.60 166.46 40.13 33.23 22.55 14.72
[…]
BiLLM 6 9537 18405.85 28123.58 71.65 42.41 30.20 18.65
[…]
4.2 Main Results
[…]
Table 2: Accuracy of PIQA, ARC-e, ARC-c, HellaSwag, Winogrande and average accuracy of all datasets with BiLLM and our methods on OPT and LLaMA family.
Model Method Activation Bits PIQA ARC-e Arc-c HellaSwag Winogrande Avg.
- 16 76.33 65.61 30.55 50.51 65.35 57.67
BiLLM 16 59.63 36.83 17.06 30.14 51.30 38.99
DBellQuant 16 70.29 55.76 24.72 37.81 58.71 49.46
[…]
DBellQuant 8 69.10 54.63 24.91 38.19 57.38 48.84
[…]
DBellQuant 6 68.12 52.44 23.38 37.04 57.30 47.65
- 16 78.40 67.34 38.14 56.45 67.01 61.46
BiLLM 16 61.92 38.93 21.58 32.78 53.67 41.77
DBellQuant 16 67.74 49.37 24.23 39.55 58.80 47.94
[…]
DBellQuant 8 67.41 47.31 26.19 38.89 58.80 47.72
[…]
DBellQuant 6 65.29 46.46 25.77 37.28 54.85 45.94
- 16 78.40 69.28 40.02 56.69 67.25 62.32
BiLLM 16 60.34 36.87 21.59 30.24 51.62 40.13
DBellQuant 16 63.98 42.85 23.89 34.82 56.27 44.36
[…]
DBellQuant 8 62.56 42.42 23.03 34.44 54.38 43.37
[…]
DBellQuant 6 61.10 37.5 22.18 31.94 53.85 41.32
- 16 79.65 80.09 50.51 60.18 72.77 68.64
BiLLM 16 57.51 33.75 18.52 31.63 53.12 38.90
DBellQuant 16 62.35 44.11 19.97 33.19 55.96 43.12
[…]
DBellQuant 8 61.70 40.95 18.40 32.95 55.09 41.82
[…]
DBellQuant 6 58.00 37.63 18.77 31.83 51.92 39.64
4.3 Ablation Experiments
Optimization Objective
In the definition of DTNP, we employed the L1 loss as described in Eq. 6. To evaluate the effectiveness of the L1 loss, we tested by replacing it with L2 loss and do the same experiments. We observed that the overall performance was inferior to that achieved by L1 loss. Therefore, we selected L1 loss as the loss function for DTNP. Additionally, during the activation-aware initialization process, the parameter $\epsilon$ is introduced as a hyperparameter. We conducted experiments with different values of $\epsilon$, and the results demonstrated that all tested values of $\epsilon$ consistently outperformed prior algorithms. For LLaMA-2-7B, setting $\epsilon$ to 0.85 yielded the best performance. Results are shown in Tab. 4.
Impact of Block Size
We investigated the impact of block size on the quantization performance of DBellQuant by experimenting with block sizes ranging from 64 to 256 columns. The results are shown in Tab. 3. Consistent with other PTQ methods, smaller block sizes yield lower perplexity, but it will increase the diversity of quantization results also increase the weighting overhead. In all experiments shown before we set the block size to 128 because we think it can better balance the bit-width and quantization effect. Furthermore, our method consistently outperforms the current baseline approaches across various block sizes, demonstrating its robustness and generalizability under different configurations.
Figure 6: Top: Visualization of single-bell weights distribution from different blocks of different layers before applying DBellQuant. Bottom: Visualization of dual-bell weights distribution from different blocks of different layers after applying DBellQuant.
Activation Bit-width Comparisons
We compare the results across different models with different activation bit-widths shown in Fig. 7. Specifically, we observe that when activations are quantized to 8 bits, the model’s perplexity remains nearly unchanged and, in some cases, even decreases. This demonstrates that our approach alleviates the challenges associated with activation quantization. Furthermore, we found that even under lower-bit quantization, such as 6-bit, our model is able to largely maintain its original performance, with only a minimal increase in perplexity. It is noteworthy that for large-scale models such as LLaMA-2 13B and 70B, the perplexity degradation is almost negligible, highlighting the effectiveness and efficiency of our approach in models with substantial parameters.
Method Avtivation Bits Block Size WikiText2 $\downarrow$ C4 $\downarrow$
BiLLM 16 64 20.12 24.46
DBellQuant 8 64 13.67 16.99
DBellQuant 6 64 15.65 19.71
[…]
DBellQuant 8 128 18.65 23.02
DBellQuant 6 128 21.69 25.91
[…]
DBellQuant 8 256 22.34 23.37
DBellQuant 6 256 24.68 28.52
Table 3: Performance of different block size.
Figure 7: Performance of different activation bit-widths.
Method Loss Function Avtivation Bits WikiText2 $\downarrow$ C4 $\downarrow$
DBellQuant L2 8 18.90 24.11
DBellQuant L1 8 18.65 23.02
[…]
DBellQuant L1 6 21.69 25.91
(a) Performance of different loss function.
Method $\epsilon$ WikiText2 $\downarrow$ C4 $\downarrow$
DBellQuant 0.75 19.57 22.92
DBellQuant 0.8 18.66 22.81
DBellQuant 0.85 17.91 21.83
DBellQuant 0.9 18.29 22.52
(b) Performance of different $\epsilon$
Table 4: Ablation study on LLaMA-2-7B,results are measured by perplexity.
5 Conclusion
In this work, we propose DBellQuant, an efficient PTQ method that enables simultaneous weight binarization and activation quantization for LLMs. By analyzing weight distributions conducive to binarization, we design a novel and efficient Learnable Transformation for Dual-Bell algorithm. Specifically, we introduce two customized loss functions and an early stopping mechanism to achieve the desired dual-bell transformation. Furthermore, we analyze and demonstrate how this transformation improves activation quantization, allowing us to achieve near 1-bit weight compression and 6-bit activation quantization with minimal performance degradation—marking the first time such results have been achieved under a PTQ framework. Experiments on open-source LLM families show that DBellQuant significantly advances the performance frontier of SOTA binary PTQ methods.
Appendix A Appendix
A.1 Limitations
Currently, our work only supports quantizing activations to 6 bits. When attempting to quantize activations to lower bit-widths, the model collapses, resulting in a significant drop in performance. However, some recent studies have demonstrated the ability to quantize activations to 4 bits while maintaining competitive model performance. Inspired by these advancements, we aim to adopt similar approaches to further optimize our models. Specifically, we seek to not only binarize weights but also quantize activations to even lower bit-widths, enabling easier deployment and faster computation.
A.2 Broader Impacts
Our work demonstrates the feasibility of simultaneously binarizing weights and quantizing activations to 6 bits in LLMs while maintaining competitive performance. This approach significantly reduces the computational and memory overhead associated with LLM deployment, making them more accessible for resource-constrained environments. By enabling efficient inference, our method contributes to the democratization of advanced AI technologies, reducing the environmental impact of large-scale model deployment. Furthermore, it opens new avenues for research in ultra-low-bit quantization, fostering innovation in model efficiency and scalability.
A.3 Proof for Theorem. 1
Proof.
Problem Setup: By assumption, the rows of the original weight matrix $\mathbf{W}$ are sampled independently from Gaussian distributions:
$\mathbf{w}_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2}),\quad\text{where }\mu_{i}\in\mathbb{R}\text{ and }\sigma_{i}>0\text{ for all }i.$
We aim to learn a transformation matrix $\mathbf{T}\in\mathbb{R}^{m\times m}$ such that the rows of the resulting matrix $\mathbf{W}^{\prime}=\mathbf{W}\mathbf{T}$ follow a bimodal distribution.
Learnable Transformation Definition: The transformed matrix is defined as:
$\mathbf{W}^{\prime}=\mathbf{W}\mathbf{T},$
where $\mathbf{T}$ is a learnable matrix that modulates the distribution of each row $\mathbf{w}_{i}^{\prime}$ of $\mathbf{W}^{\prime}$. Since the rows of $\mathbf{W}$ are Gaussian-distributed, the linear transformation by $\mathbf{T}$ initially results in a new Gaussian distribution for each row:
$\mathbf{w}_{i}^{\prime}\sim\mathcal{N}(\mu_{i}^{\prime},\sigma_{i}^{\prime 2}),$
where $\mu_{i}^{\prime}=\mu_{i}\mathbf{T}$ and $\sigma_{i}^{\prime 2}=\mathbf{T}^{\top}\Sigma_{i}\mathbf{T}$, with $\Sigma_{i}=\mathrm{diag}(\sigma_{i}^{2})$ being the covariance of $\mathbf{w}_{i}$.
Inducing a Bimodal Distribution: To map the Gaussian-distributed rows $\mathbf{w}_{i}^{\prime}$ into a bimodal distribution, we note that a bimodal distribution can be expressed as a Gaussian mixture model:
$g(x)=\pi\mathcal{N}(x;\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(x;\mu_{2},\sigma_{2}^{2}),$
where $\pi\in(0,1)$ is the mixing coefficient, and $\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2}$ are the parameters of the mixture components. To achieve this, $\mathbf{T}$ is learned to ensure that the linear transformation $\mathbf{W}\mathbf{T}$ reshapes the original Gaussian distribution into a mixture of two Gaussians.
Parameter Optimization: The learnable matrix $\mathbf{T}$ is optimized using a loss function $L$ that minimizes the Kullback-Leibler (KL) divergence between the empirical distribution of the rows of $\mathbf{W}^{\prime}$ and the target bimodal distribution:
$L=\mathrm{KL}\left(p(\mathbf{w}_{i}^{\prime})\,\|\,\pi\mathcal{N}(\mu_{1},\sigma_{1}^{2})+(1-\pi)\mathcal{N}(\mu_{2},\sigma_{2}^{2})\right).$
The optimization process adjusts the entries of $\mathbf{T}$ to align the transformed rows $\mathbf{w}_{i}^{\prime}$ with the desired bimodal distribution.
Conclusion: The existence of such a learnable matrix $\mathbf{T}$ ensures that the rows of the transformed matrix $\mathbf{W}^{\prime}=\mathbf{W}\mathbf{T}$ can follow a bimodal distribution. This completes the proof. ∎
A.4 Analysis of the reasons doubel-bell distribution more suitable for binarization compared to a single-bell distribution.
Directly proving that a dual-bell distribution is more suitable for binarization compared to a single-bell distribution can be challenging, as it requires setting numerous additional conditions. However, this problem becomes significantly simpler when approached from the perspective of value adjustments. By reducing the magnitude of larger absolute values and increasing smaller absolute values in a bimodal distribution, all values can be shifted closer to two central points, effectively creating a double-bell-like distribution. We can demonstrate that this approach reduces the quantization loss introduced by binarization, thereby supporting the suitability of double-bell distributions for this purpose.
Theorem 2.
Given an input calibration activation $x\in\mathbb{R}^{n\times 1}$ and a weight vector $\boldsymbol{w}\in\mathbb{R}^{n\times 1}$, where $w_{i}$ is extracted from the weight matrix $\boldsymbol{W}\in\mathbb{R}^{n\times n}$ along a specific channel, we define the weight vector $\boldsymbol{w}$ as the union of two sets: - A set of several outliers with large absolute values, denoted as $U_{o}=\{o_{1}^{*},o_{2}^{*},\dots,o_{k}^{*}\}$, where $|o_{i}^{*}|\gg 0$ for $i\in\{1,\dots,k\}$; - A set of normal values with small absolute values, denoted as $U_{n}=\{n_{1},n_{2},\dots,n_{n-k}\}$, where $|n_{j}|\approx 0$ for $j\in\{1,\dots,n-k\}$. $\boldsymbol{w}=U_{o}\cup U_{n}.$ We now define a new weight vector $\boldsymbol{w}_{\text{new}}$ as follows:
  • $\boldsymbol{w}_{\text{new}}=[n_{1},n_{2},\dots,\gamma o_{1}^{*},\gamma o_{2}^{*},\dots,\gamma o_{k}^{*},\dots,n_{n-k}]$, where $\gamma\in(\frac{1}{2},1)$.
  • Alternatively, $\boldsymbol{w}_{\text{new}}=[\eta n_{1},\eta n_{2},\dots,o_{1}^{*},o_{2}^{*},\dots,o_{k}^{*},\dots,\eta n_{n-k}]$, where $\eta\in(1,2)$.
Then, the quantization error induced by $\boldsymbol{w}_{\text{new}}$, defined as $\|x\cdot\boldsymbol{w}-x\cdot\text{binarized}(\boldsymbol{w}_{\text{new}})\|$, is strictly smaller than the original quantization error $\|x\cdot\boldsymbol{w}-x\cdot\text{binarized}(\boldsymbol{w})\|$ in both cases.
Proof.
A.4.1 Scaling Up Small Values Reduces Quantization Error
Consider the scenario where the input vector is $X=[2,2,2,\dots,2]_{n}$. Assume the weights in a single channel, $W$, are given by $[\alpha_{1},\alpha_{2},...,\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $[\alpha_{i}\in U_{1}]$ is the set of values with very large absolute magnitudes, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $[\beta_{i}\in U_{2}]$ are values with magnitudes close to zero, satisfying $\sum_{i=1}^{n-k}\beta_{i}=0$. $W=U_{1}\cup U_{2}$. This structure is common in practice, as weight distributions in neural networks often exhibit a few dominant values and many small ones. As we observe, the distribution of weights along the channel dimension is mostly not symmetric around zero. Instead, it tends to be biased, with the majority leaning either towards positive or negative values. Consequently, the extreme values are predominantly either entirely positive or entirely negative, so we assume $\alpha_{i}>0$.
The product of the input $X$ and the weight vector $W$ is:
$X\cdot W^{T}=2(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k})=2(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})=2A$ (7)
since the sum of the $\beta_{i}$ is zero.
According to the quantization function, the mean value $M$ is:
$M=\frac{\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}}{n}=\frac{A}{n}$ (8)
The absolute mean value, $AbsMean$, is defined as:
$\begin{split}AbsMean&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|\beta_{1}-\frac{A}{n}|+|\beta_{2}-\frac{A}{n}|+\dots+|\beta_{n-k}-\frac{A}{n}|}{n}\end{split}$ (9)
Given that $[\beta_{1},\beta_{2},\dots,\beta_{n-1}]$ are all very small in magnitude compared to $\frac{\alpha}{n}$, we can simplify the above as:
$\begin{split}AbsMean&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-\beta_{1})+(\frac{A}{n}-\beta_{2})+\dots+(\frac{A}{n}-\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{2kA}{n})}{n}\end{split}$ (10)
Here, the sum of the $\beta_{i}$ vanishes due to their zero sum constraint.
Therefore, the dequantized value for $\alpha$ is:
$AbsMean+M=\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (11)
and for each $\beta_{i}$:
$-AbsMean+M=-\frac{2(A-\frac{2kA}{n})}{n}+\frac{A}{n}$ (12)
To reduce the quantization error associated with the small-magnitude weights, we can scale them up by a factor $m>1$, while correspondingly scaling down the input. Specifically, we multiply each $\beta_{i}$ by $m$ and divide the associated input elements by $m$. The new input becomes $X_{new}=[2,2,\dots,2,\frac{2}{m},\frac{2}{m},\dots,\frac{2}{m}]_{n}$, and the new weights are $W_{new}=[\alpha_{1},\alpha_{2},...,\alpha_{k},m\beta_{1},m\beta_{2},\dots,m\beta_{n-k}]$.
The output remains unchanged:
$X_{new}\cdot W_{new}^{T}=2(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})+\frac{2}{m}\cdot m\beta_{1}+\frac{2}{m}\cdot m\beta_{2}+\dots+\frac{2}{m}\cdot m\beta_{n-k}=2A$ (13)
This invariance is crucial: the scaling operation does not affect the original computation, but it can impact the quantization error.
For the new weights, the mean value is:
$M_{new}=\frac{\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+m\beta_{1}+m\beta_{2}+\dots+m\beta_{n-k}}{n}=\frac{A+m(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}=\frac{A}{n}$ (14)
since the sum of the $\beta_{i}$ is zero.
The new absolute mean value is:
$\begin{split}AbsMean_{new}&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|m\beta_{1}-\frac{\alpha}{n}|+|m\beta_{2}-\frac{\alpha}{n}|+\dots+|m\beta_{n-k}-\frac{\alpha}{n}|}{n}\end{split}$ (15)
If $m$ is chosen such that $\frac{\alpha}{n}$ remains larger than all $m\beta_{i}$, the simplification proceeds as before:
$\begin{split}AbsMean_{new}&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{A}{n}-m\beta_{1})+(\frac{A}{n}-m\beta_{2})+\dots+(\frac{A}{n}-m\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}-m(\beta_{1}+\beta_{2}+\dots+\beta_{n-k})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{kA}{n})}{n}\end{split}$ (16)
Thus, $AbsMean_{new}$ and $M_{new}$ are identical to $AbsMean$ and $M$, and the dequantized values are unchanged. This demonstrates that scaling up the small weights does not affect the mean or absolute mean, but it can improve the quantization error, as we analyze next.
Let us now analyze the quantization error. The original output is $2\alpha$. For convenience, let $N=(n-k)(-AbsMean+M)$. The quantized output is a sum of the dequantized values for all weights.
$A>0$ Both before and after scaling, the quantization output contains the term:
$2k(AbsMean+M)=2k\left(\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}\right)$ (17)
For $n$ typically greater than 100 and $k\ll n$ it holds that:
$0<2k(AbsMean+M)<2A$ (18)
This is because the quantized value is always less than the original due to the averaging effect.
For the term $-AbsMean+M$:
$\begin{split}-AbsMean+M&=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}\\ &=-\frac{A}{n}+\frac{2kA}{n^{2}}\\ &=\frac{A}{n}\left(\frac{2k}{n}-1\right)\end{split}$ (19)
Since $n>100$ ,$A>0$ and $k\ll n$ this value is negative and its magnitude is small.
The quantization output before scaling is $2k(AbsMean+M)+2N$, and after scaling is $2k(AbsMean+M)+N$. Because $N<0$ and $2(AbsMean+M)<2A$, we have:
$2k(AbsMean+M)+2N<2k(AbsMean+M)+N<2A$ (20)
This shows that scaling up the small weights reduces the quantization error, as the quantized output moves closer to the original value.
if $\alpha_{i}<0$, the proof is similar.
In summary, scaling up the small weights (and correspondingly scaling down the input) does not change the original computation, but it systematically reduces the quantization error by making the quantized output more faithful to the original.
A.4.2 Scaling Down Large Values Reduces Quantization Error
Now, let us consider the scenario where we scale down the large-magnitude weight. Let the input $X=[1,1,1,\dots,1]_{n}$, and the weights $W=[m\alpha_{1},m\alpha_{2},...,m\alpha_{k},\beta_{1},\beta_{2},\beta_{3},\dots,\beta_{n-k}]$, where $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are large values, satisfying $\sum_{i=1}^{k}\alpha_{i}=A$ and $\alpha_{i}>0$, $[\beta_{1},\dots,\beta_{n-k}]$ are small, and $\sum_{i=1}^{n-k}\beta_{i}=0$.
The output is:
$X\cdot W^{T}=m(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}=mA$ (21)
The mean value is:
$M=\frac{m\alpha_{1}+m\alpha_{2}+\dots+m\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}}{n}=\frac{mA}{n}$ (22)
The absolute mean is:
$\begin{split}AbsMean&=\frac{|m\alpha_{1}-\frac{mA}{n}|+|m\alpha_{2}-\frac{mA}{n}|+\dots+|m\alpha_{k}-\frac{mA}{n}|+|\beta_{1}-\frac{mA}{n}|+|\beta_{2}-\frac{mA}{n}|+\dots+|\beta_{n-k}-\frac{mA}{n}|}{n}\end{split}$ (23)
Since $\frac{mA}{n}$ is much larger than the $\beta_{i}$, we can simplify:
$\begin{split}AbsMean&=\frac{(m\alpha_{1}-\frac{mA}{n})+(m\alpha_{2}-\frac{mA}{n})+\dots+(m\alpha_{k}-\frac{mA}{n})+(\frac{mA}{n}-\beta_{1})+(\frac{mA}{n}-\beta_{2})+\dots+(\frac{mA}{n}-\beta_{n-k})}{n}\\ &=\frac{m(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})+(n-2k)\frac{mA}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-1})}{n}\\ &=\frac{mA+(n-2k)\frac{mA}{n}}{n}\\ &=\frac{2m(A-\frac{kA}{n})}{n}\end{split}$ (24)
The dequantized value for $m\alpha$ is:
$AbsMean+M=\frac{2m(A-\frac{kA}{n})}{n}+\frac{mA}{n}=\frac{3mA-\frac{2mkA}{n}}{n}$ (25)
and for each $\beta_{i}$:
$-AbsMean+M=-\frac{2m(A-\frac{A}{n})}{n}+\frac{mA}{n}=\frac{-mA+\frac{2mkA}{n}}{n}$ (26)
To scale down the large value $m\alpha$, we divide it by $m$ ($m>1$) and multiply the corresponding input element by $m$. The new input is $X_{new}=[m,1,1,\dots,1]_{n}$, and the new weights are $W_{new}=[\alpha_{1},\alpha_{2},\dots,\alpha_{k},\beta_{1},\beta_{2},\dots,\beta_{n-k}]$.
The output remains unchanged:
$X_{new}\cdot W_{new}^{T}=m(\alpha_{1}+\alpha_{2}+\dots+\alpha_{k})+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}=mA$ (27)
For the new weights, the mean is:
$M_{new}=\frac{\alpha_{1}+\alpha_{2}+\dots+\alpha_{k}+\beta_{1}+\beta_{2}+\dots+\beta_{n-k}}{n}=\frac{A}{n}$ (28)
The new absolute mean is:
$\begin{split}AbsMean_{new}&=\frac{|\alpha_{1}-\frac{A}{n}|+|\alpha_{2}-\frac{A}{n}|+\dots+|\alpha_{k}-\frac{A}{n}|+|\beta_{1}-\frac{A}{n}|+|\beta_{2}-\frac{A}{n}|+\dots+|\beta_{n-1}-\frac{A}{n}|}{n}\end{split}$ (29)
With appropriate $m$, we have:
$\begin{split}AbsMean_{new}&=\frac{(\alpha_{1}-\frac{A}{n})+(\alpha_{2}-\frac{A}{n})+\dots+(\alpha_{k}-\frac{A}{n})+(\frac{\alpha}{n}-\beta_{1})+(\frac{\alpha}{n}-\beta_{2})+\dots+(\frac{\alpha}{n}-\beta_{n-k})}{n}\\ &=\frac{(\alpha_{1}+\alpha_{2}++\dots+\alpha_{k})+(n-2k)\frac{A}{n}-(\beta_{1}+\beta_{2}+\dots+\beta_{n-1})}{n}\\ &=\frac{A+(n-2k)\frac{A}{n}}{n}\\ &=\frac{2(A-\frac{kA}{n})}{n}\end{split}$ (30)
The dequantized value for $\alpha_{i}$ is:
$AbsMean_{new}+M_{new}=\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{3A-\frac{2kA}{n}}{n}$ (31)
and for each $\beta_{i}$:
$-AbsMean_{new}+M_{new}=-\frac{2(A-\frac{kA}{n})}{n}+\frac{A}{n}=\frac{-A+\frac{2kA}{n}}{n}$ (32)
Let us now examine the quantization error. The original output is $m\alpha$. The quantization output before scaling is:
$\begin{split}&k(AbsMean+M)+(n-k)(-AbsMean+M)\\ &=k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-mA+\frac{2mkA}{n}}{n}\end{split}$ (33)
The quantization output after scaling is:
$\begin{split}&km(AbsMean_{new}+M_{new})+(n-k)(-AbsMean_{new}+M_{new})\\ &=km\frac{3A-\frac{2kA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &=k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\end{split}$ (34)
Because $A>0$, for $n>100$ ,$\alpha>0$ and $k\ll n$ it holds that:
$0<k\frac{3mA-\frac{2mkA}{n}}{n}<mA$ (35)
and
$\frac{-A+\frac{2kA}{n}}{n}=\frac{A}{n}\left(\frac{2k}{n}-1\right)<0$ (36)
Comparing the quantization outputs, we see:
$\begin{split}&k\frac{3mA-\frac{2mkA}{n}}{n}+m(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &<k\frac{3mA-\frac{2mkA}{n}}{n}+(n-k)\frac{-A+\frac{2kA}{n}}{n}\\ &<mA\end{split}$ (37)
If $\alpha_{i}<0$, the proof is similar.
A.5 Perplexity on C4 dataset
As shown in Tab.5, we compare the perplexity of the OPT and LLaMA families across different model sizes on C4 dataset.
[…]
Method Activation Bits OPT-1.3B OPT-2.7B OPT-6.7B LLaMA-1-7B LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
Full Precision 16 16.07 14.34 12.71 7.34 07.26 06.73 5.71
RTN 16 $9$.56 23492.89 9617.07 194607.78 115058.76 46250.21 314504.09
GPTQ 16 6364.65 6703.36 5576.82 186229.5 67954.04 19303.51 13036.32
PB-LLM 16 168.12 222.15 104.78 76.63 80.69 184.67 NAN
BiLLM 16 64.14 44.77 42.13 46.96 39.38 25.87 17.30
[…]
BiLLM 8 74.56 61.99 40.91 47.13 40.91 21.45 17.72
[…]
BiLLM 6 7348 13445.21 63.41 61.65 63.41 37.66 19.43
[…]
A.6 Visualization of distribution of activation before and after DBellQuant
In this section, we present the changes in the distribution of activation values before and after applying DBellQuant as shown in Fig.8. It is evident that the extreme values in the activation have been significantly reduced by a factor of 5 to 10; for instance, the maximum value decreases from approximately 3 to around 0.4. Previous studies have highlighted that one of the primary challenges in low-bit quantization of activations lies in the presence of large outliers, which expand the activation range and, consequently, amplify quantization errors. By applying DBellQuant, the activation range is effectively compressed from [-3, 3] to [-0.4, 0.4], dramatically alleviating the difficulty of quantization. This reduction in range establishes highly favorable conditions for further exploration of lower-bit quantization, such as 8-bit or even 6-bit implementations.
Figure 8: Visualization results of distribution of activation before and after DBellQuant across different blocks.